### 02 Dec 2023 v1

# On the Moser’s Bernstein Theorem

Abstract.
In this paper, we prove the following version of the famous Bernstein’s theorem: Let $X\subset\mathbb{R}^{n+k}$ be a closed and connected set with Hausdorff dimension $n$. Assume that $X$ satisfies the monotonicity formula at $p\in X$. Then, the following statements are equivalent:
  • (1)
    $X$ is an affine linear subspace;
  • (2)
    $X$ is a definable set that is Lipschitz regular at infinity and its geometric tangent cone at infinity, $C(X,\infty)$, is a linear subspace;
  • (3)
    $X$ is a definable set, blow-spherical regular at infinity and $C(X,\infty)$ is a linear subspace;
  • (4)
    $X$ is an LNE at infinity definable set and $C(X,\infty)$ is a linear subspace;
  • (5)
    the density of $X$ at infinity is 1.
As a consequence, we obtain a proof of the following generalization of Bernstein’s theorem: Let $X\subset\mathbb{R}^{n+1}$ be a closed and connected set with Hasdorff dimension $n$. Assume that $X$ satisfies the monotonicity formula at $p\in X$ and there are compact sets $K\subset\mathbb{R}^{n}$ and $\tilde{K}\subset\mathbb{R}^{n+1}$ such that $X\setminus\tilde{K}$ is a minimal hypersurface that is the graph of a $C^{2}$ smooth function $u\colon\mathbb{R}^{n}\setminus K\to\mathbb{R}$. Assume that $u$ has bounded derivative whenever $n>7$. Then $X$ is a hyperplane. Several other consequences are also presented.
Key words and phrases:
Bernstein Theorem, Moser’s Bernstein Theorem, Minimal surfaces, monotonicity formula, Lipschitz regularity
2010 Mathematics Subject Classification:
53A10; 53C42; 14J17; 14P10
The first named author was partially supported by CNPq-Brazil grant 310438/2021-7. This work was supported by the Serrapilheira Institute (grant number Serra – R-2110-39576).
1. Introduction
We start this article by reminding the famous Bernstein’s problem:
Bernstein’s problem.
If the graph of a function on $\mathbb{R}^{n}$ is a minimal hypersurface in $\mathbb{R}^{n+1}$, does this imply that the function is linear?
This problem is false for $n$ at least 8 (see [7]), but the answer was proved to be positive in the cases $n=2$ by Bernstein, $n=3$ by DeGiorgi, $n=4$ by Almgren and $n\leq 7$ by Simons. We can summarize the positive answers as follows:
Theorem 1.1.
Let $M\subset\mathbb{R}^{n+1}$ be a minimal hypersurface which is a graph of $C^{2}$ smooth function $u\colon\mathbb{R}^{n}\to\mathbb{R}$ with $2\leq n\leq 7$. Then $u$ is a linear function and $M$ must be a hyperplane.
Furthermore, the celebrated theorem due to J. Moser in [27], called Moser’s Bernstein Theorem, says the following:
Theorem 1.2 (Moser’s Bernstein Theorem).
Let $M\subset\mathbb{R}^{n+1}$ be a minimal hypersurface which is a graph of an entire Lipschitz function $u\colon\mathbb{R}^{n}\to\mathbb{R}$. Then $u$ is a linear function and $M$ must be a hyperplane.
Let us remark that, in general, Moser’s Bernstein Theorem does not hold in higher codimension. Indeed, Lawson and Osserman in [24] presented a minimal cone which is the graph of a Lipschitz mapping, but is not an affine linear subspace, more precisely, they presented the following:
Example 1.3 (Theorem 7.1 in [24]).
The graph of the Lipschitz mapping $f\colon\mathbb{R}^{4}\to\mathbb{R}^{3}$ given by $f(0)=0$ and
$f(x)=\frac{\sqrt{5}}{2}\|x\|\eta\left(\frac{x}{\|x\|}\right),\quad\forall x\not=0,$
is a minimal cone, where $\eta\colon\mathbb{S}^{3}\to\mathbb{S}^{2}$ is the Hopf mapping given by
$\eta(z_{1},z_{2})=(|z_{1}|^{2}-|z_{2}|^{2},2z_{1}\bar{z}_{2}).$
Several other mathematicians approached Moser’s Bernstein Theorem in higher codimension (for example, see [19], [20], [21] and [47]).
In this direction, we prove here the following result:
Theorem 5.1.
Let $X\subset\mathbb{R}^{n+k}$ be a connected closed set with Hausdorff dimension $n$. Then, $X$ is a $n$-dimensional affine linear subspace if and only if we have the following:
  1. (1)
    $X$ satisfies the monotonicity formula at $p\in X$;
  2. (2)
    there are compact sets $K\subset\mathbb{R}^{n}$ and $\tilde{K}\subset\mathbb{R}^{n+k}$ such that $X\setminus\tilde{K}$ is the graph of a $C^{1}$ smooth function $u\colon\mathbb{R}^{n}\setminus K\to\mathbb{R}^{k}$ with bounded derivative;
  3. (3)
    $\mathcal{N}(X,\infty)$ is a linear subspace.
In order to know, for a subset $X\subset\mathbb{R}^{n+k}$, with Hausdorff dimension $n$ and $p\in X$, we say that $X$ satisfies the monotonicity formula at $p$ if the function $\theta^{n}(X,p,\cdot)\colon(0,+\infty)\to[1,+\infty)$ given by
$\theta^{n}(X,p,r)=\frac{\mathcal{H}^{n}(X\cap B_{r}^{n}(p))}{\mu_{n}r^{n}}$
is well-defined and non-decreasing and, moreover, satisfies the following:
  • (i)
    $X$ is $C^{1}$-smooth at $p$ if and only if $\theta^{n}(X,p):=\lim\limits_{r\to 0^{+}}\theta^{n}(X,p,r)=1$;
  • (ii)
    $\theta^{n}(X,p,\cdot)$ is a constant function if and only if $X$ is a cone with vertex at $p$,
where $\mu_{n}$ is the volume of the $n$-dimensional Euclidean unit ball, $\mathcal{H}^{n}(A)$ is the $n$-dimensional Hausdorff measure of $A$ and $B_{r}^{m}(p)\subset\mathbb{R}^{m}$ is the open Euclidean ball center at $p$ of radius $r>0$. Moreover, $\mathcal{N}(X,\infty)$ is the union of all hyperplanes $T$ such that there is a sequence $\{x_{j}\}_{j}\subset X\setminus{\rm Sing}_{1}(X)$ such that $\lim\|x_{j}\|=+\infty$ and $T_{x_{j}}X$ converges to $T$, where ${\rm Sing}_{1}(X)$ is the set of points $x\in X$ such that $X\cap U$ is not a $C^{1}$ smooth submanifold of $\mathbb{R}^{n+k}$, for any open neighbourhood $U\subset\mathbb{R}^{n+k}$ of $x$.
Remark 1.4.
If $X\subset\mathbb{R}^{n+k}$ is a complete minimal submanifold or a closed area-minimizing set, then $X$ satisfies the monotonicity formula at any $p\in X$. In particular, any entire pure dimensional complex analytic set $X\subset\mathbb{C}^{n}$ satisfies the monotonicity formula at any $p\in X$.
In Subsection 5.1, by presenting several examples, we show that Theorem 5.1 is sharp in the sense that no hypothesis of it can be dropped.
As consequences of Theorem 5.1, we obtain the following generalizations of Theorems 1.1 and 1.2:
Corollary 5.7.
Let $X\subset\mathbb{R}^{n+1}$ be a connected closed set with Hausdorff dimension $n$ and that satisfies the monotonicity formula at $p\in X$. Suppose $2\leq n\leq 7$ and that there are compact sets $K\subset\mathbb{R}^{n}$ and $\tilde{K}\subset\mathbb{R}^{n+1}$ such that $X\setminus\tilde{K}$ is a minimal hypersurface that is the graph of a $C^{2}$ smooth function $u\colon\mathbb{R}^{n}\setminus K\to\mathbb{R}$. Then $u$ is the restriction of an affine function and, in particular, $X$ is a hyperplane.
Corollary 5.6.
Let $X\subset\mathbb{R}^{n+1}$ be a connected closed set with Hausdorff dimension $n$ and that satisfies the monotonicity formula at $p\in X$. Suppose that there are compact sets $K\subset\mathbb{R}^{n}$ and $\tilde{K}\subset\mathbb{R}^{n+1}$ such that $X\setminus\tilde{K}$ is a minimal hypersurface that is the graph of a $C^{2}$ smooth function $u\colon\mathbb{R}^{n}\setminus K\to\mathbb{R}$ with bounded derivative. Then $u$ is the restriction of an affine function and, in particular, $X$ is a hyperplane.
Another natural question is to ask whether Moser’s Bernstein Theorem can be generalized to the parametric case. In other words, it is natural to ask the following:
Question 1.
Given a smooth minimal hypersurface $M\subset\mathbb{R}^{n+1}$ which is bi-Lipschitz homeomorphic to $\mathbb{R}^{n}$, is $M$ an affine linear subspace?
Recently, in [14] and [38] approached Question 1 in higher codimension for complex analytic sets. For instance, it was proved that a pure dimensional complex analytic set which is Lipschitz regular at infinity (see Definition 2.6) must be an affine linear subset.
In this article, we use the main ideas of [14] and [38] to give some partial positive answers to Question 1, even for higher codimension. In particular, we prove the following:
Theorem 4.1.
Let $X\subset\mathbb{R}^{n+k}$ be a closed and connected set and $n=\dim_{H}X$. Assume that $X$ satisfies the monotonicity formula at $p\in X$. Then the following statements are equivalent:
  • (1)
    $X$ is an affine linear subspace;
  • (2)
    $X$ is a definable set, blow-spherical regular at infinity and $C(X,\infty)$ is a linear subspace;
  • (3)
    $X$ is a definable set that is Lipschitz regular at infinity and $C(X,\infty)$ is a linear subspace;
  • (4)
    $X$ is an LNE at infinity definable set and $C(X,\infty)$ is a linear subspace;
  • (5)
    $\theta^{n}(X):=\lim\limits_{r\to+\infty}\frac{\mathcal{H}^{n}(X\cap B_{r}^{n}(0))}{\mu_{n}r^{n}}=1$.
We obtain several consequences, for instance, we recover the results proved in [14] and [40].
The main tool to prove Theorem 4.1 is the Kurdyka-Raby’s formula at infinity.
Theorem 3.1.
(Kurdyka-Raby’s formula at infinity). Let $X\subset\mathbb{\mathbb{R}}^{n+k}$ be a definable set in an o-minimal structure on $\mathbb{R}$ and $n=\dim X$. Let $C_{1},...,C_{m}$ be the connected components of $C_{{\rm Smp}}(X,\infty)$. Then, $\theta^{n}(X)$ is defined and for each $x\in\mathbb{R}^{n+k}$, we have
$\theta^{n}(X)=\lim\limits_{r\to+\infty}\frac{\mathcal{H}^{n}(X\cap B_{r}^{n+k}(x))}{\mu_{n}r^{n}}=\sum\limits_{j=1}^{m}k_{X,\infty}(C_{j})\frac{\mathcal{H}^{n}(C_{j}\cap B_{1}^{n+k})}{\mu_{n}}.$
We prove this formula in Section 3. Still in Section 3, we present some consequences of Theorem 3.1. For instance, we obtain the o-minimal Chow’s Theorem proved in [31, Corollary 4.5]:
Corollary 3.9 (O-minimal Chow’s Theorem).
Let $X\subset\mathbb{C}^{n}$ be a dimensional entire complex analytic set with $\dim_{\mathbb{C}}X=d$. Then, $X$ is a complex algebraic set if and only if $X$ is definable in an o-minimal structure on $\mathbb{R}$.
2. Preliminaries
All the sets considered in this paper are supposed to be equipped with the induced Euclidean metric.
2.1. O-minimal structures
Definition 2.1.
A structure on $\mathbb{R}$ is a collection $\mathcal{S}=\{\mathcal{S}_{n}\}_{n\in\mathbb{N}}$ where each $\mathcal{S}_{n}$ is a set of subsets of $\mathbb{R}^{n}$, satisfying the following axioms:
  • 1)
    All algebraic subsets of $\mathbb{R}^{n}$ are in $\mathcal{S}_{n}$;
  • 2)
    For every $n$, $\mathcal{S}_{n}$ is a Boolean subalgebra of the powerset of $\mathbb{R}^{n}$;
  • 3)
    If $A\in\mathcal{S}_{m}$ and $B\in S_{n}$, then $A\times B\in\mathcal{S}_{m+n}$.
  • 4)
    If $\pi\colon\mathbb{R}^{n+1}\to\mathbb{R}^{n}$ is the projection on the first $n$ coordinates and $A\in\mathcal{S}_{n+1}$, then $\pi(A)\in\mathcal{S}_{n}$.
The structure $\mathcal{S}$ is said to be o-minimal if, moreover, it satisfies:
  • 5)
    The elements of $\mathcal{S}_{1}$ are precisely the finite unions of points and intervals.
The elements of $\mathcal{S}_{n}$ are called the definable subsets of $\mathbb{R}^{n}$.
Throughout the paper we fix an o-minimal structure $\mathcal{S}$ on $\mathbb{R}$.
Definition 2.2.
A map $f\colon A\subset\mathbb{R}^{n}\to\mathbb{R}^{p}$ is called definable in $\mathcal{S}$ if its graph is a elements of $\mathcal{S}_{n+m}$.
In the sequel the adjective definable means definable in $\mathcal{S}$.
2.2. Dimension of definable sets
Definition 2.3.
Let $X\subset\mathbb{R}^{n}$ be a subset and let $k$ be a positive integer. The $C^{k}$ singular set of $X$, denoted by ${\rm Sing}_{k}(X)$, is the set of points $x\in X$ such that $U\cap X$ is not a smooth submanifold of $\mathbb{R}^{n}$ for any open neighbourhood $U$ of $x$. A point of ${\rm Sing}_{k}(X)$ is called a $C^{k}$ singular point (or a $C^{k}$ singularity) of $X$. If $p\in{\rm Reg}_{k}(X):=X\setminus{\rm Sing}_{k}(X)$, we say that $X$ is smooth at $p$.
Thus, if $p\in{\rm Reg}_{k}(X)=X\setminus{\rm Sing}_{k}(X)$, there is open neighbourhood $U\subset\mathbb{R}^{n}$ of $p$ such that $X\cap U$ is a $C^{k}$ smooth submanifold of $\mathbb{R}^{n}$ and, then, we define the real dimension of $X$ at $p$ by $\dim_{p}X=\dim X\cap U$. Thus, we define the dimension of $X$ by
$\dim X=\max\limits_{p\in{\rm Reg}_{1}(X)}{\dim}_{p}X.$
We say that $X$ has pure dimension, if $\dim X=\dim_{p}X$ for all $p\in{\rm Reg}_{1}(X)$.
In the case that $X\subset\mathbb{R}^{n}$ is a definable set, we have that ${\rm Sing}_{k}(X)$ is also a definable set with $\dim{\rm Sing}(X)<\dim X$ and $\dim X$ is equal to the Hausdorff dimension of $X$.
In the case that $X\subset\mathbb{C}^{n}\equiv\mathbb{R}^{2n}$ is a complex analytic set, we define the complex dimension of $X$ by $\dim_{\mathbb{C}}X=\frac{1}{2}\dim X$.
2.3. Lipschitz regularity at infinity
Definition 2.4.
Let $X\subset\mathbb{R}^{n}$ and $Y\subset\mathbb{R}^{m}$ be two subsets. We say that a mapping $\phi\colon X\rightarrow Y$ is Lipschitz if there is a constant $C\geq 0$ such that
$\|\phi(x)-\phi(y)\|\leq C\|x-y\|,\quad\forall x,y\in X.$
In this case, we also say that $\phi$ is $C$-Lipschitz. We say that a mapping $\phi\colon X\rightarrow Y$ is a lipeomorphism if $\phi$ is a homeomorphism such that $\phi$ and $\phi^{-1}$ are Lipschitz. When there is a lipeomorphism $\phi\colon X\rightarrow Y$, we say that $X$ and $Y$ are lipeomorphic.
Definition 2.5.
Let $X\subset\mathbb{R}^{n}$ and $Y\subset\mathbb{R}^{m}$ be two subsets. We say that $X$ and $Y$ are lipeomorphic at infinity, if there exist compact subsets $K\subset\mathbb{R}^{n}$ and $\widetilde{K}\subset\mathbb{R}^{m}$ such that $X\setminus K$ and $Y\setminus\widetilde{K}$ are lipeomorphic.
Definition 2.6.
A subset $X\subset\mathbb{R}^{n}$ is called Lipschitz regular at infinity if $X$ and $\mathbb{R}^{k}$ are bi-Lipschitz homeomorphic at infinity, for some $k\in\mathbb{N}$.
2.4. Lipschitz normal embedding at infinity
Let us remind the definition of inner distance. Given a path connected subset $X\subset\mathbb{R}^{m}$ the inner distance on $X$ is defined as follows: given two points $x_{1},x_{2}\in X$, $d_{X}(x_{1},x_{2})$ is the infimum of the lengths of paths on $X$ connecting $x_{1}$ to $x_{2}$. As we said in the beginning of Section 2 all the sets considered in this paper are supposed to be equipped with the induced Euclidean metric. Whenever we consider the inner distance, we emphasize it clearly.
Definition 2.7 (See [3]).
A subset $X\subset\mathbb{R}^{n}$ is called Lipschitz normally embedded (or shortly LNE) if there exists $C>0$ such that
$d_{X}(x_{1},x_{2})\leq C\|x_{1}-x_{2}\|$
for all $x_{1},x_{2}\in X$. In this case, we also say that $X$ is $C$-LNE.
Proposition 2.8 (Proposition 2.7 in [14]).
If a closed unbounded subset $X\subset\mathbb{R}^{n}$ is Lipschitz regular at infinity, then there exists a compact $K\subset\mathbb{R}^{n}$ such that each connected component of $X\setminus K$ is Lipschitz normally embedded.
In [14] was presented the following definition:
Definition 2.9.
A subset $X\subset\mathbb{R}^{n}$ is Lipschitz normally embedded at infinity (or shortly LNE at infinity) if there exists a compact subset $K\subset\mathbb{R}^{n}$ such that $X\setminus K$ is Lipschitz normally embedded.
2.5. Tangent cones
Definition 2.10.
Let $X\subset\mathbb{R}^{m}$ be an unbounded set (resp. $p\in\overline{X}$ be a non-isolated point). We say that $v\in\mathbb{R}^{m}$ is a tangent vector of $X$ at infinity (resp. $p$) if there are a sequence of points $\{x_{i}\}\subset X$ tending to infinity (resp. $p$) and a sequence of positive real numbers $\{t_{i}\}$ such that
$\lim\limits_{i\to\infty}\frac{1}{t_{i}}x_{i}=v\quad(\mbox{resp. }\lim\limits_{i\to\infty}\frac{1}{t_{i}}(x_{i}-p)=v).$
Let $C(X,\infty)$ (resp. $C(X,p)$) denote the set of all tangent vectors of $X$ at infinity (resp. $p$). We call $C(X,\infty)$ the geometric tangent cone of $X$ at infinity (resp. $p$).
Another way to present the geometric tangent cone at infinity (resp. $p$) of a subset $X\subset\mathbb{R}^{n}$ is via the spherical blow-up at infinity (resp. $p$) of $\mathbb{R}^{n}$. Let us consider the spherical blowing-up at infinity (resp. $p$) of $\mathbb{R}^{n}$, $\rho_{\infty}\colon\mathbb{S}^{n}\times(0,+\infty)\to\mathbb{R}^{n}$ (resp. $\rho_{p}\colon\mathbb{S}^{n-1}\times[0,+\infty)\to\mathbb{R}^{n}$), given by $\rho_{\infty}(x,r)=\frac{1}{r}x$ (resp. $\rho_{p}(x,r)=rx+p$).
Note that $\rho_{\infty}\colon\mathbb{S}^{n-1}\times(0,+\infty)\to\mathbb{R}^{n}\setminus\{0\}$ (resp. $\rho_{p}\colon\mathbb{S}^{n-1}\times(0,+\infty)\to\mathbb{R}^{n}\setminus\{0\}$) is a homeomorphism with inverse mapping $\rho_{\infty}^{-1}\colon\mathbb{R}^{n}\setminus\{0\}\to\mathbb{S}^{n-1}\times(0,+\infty)$ (resp. $\rho_{p}\colon\mathbb{S}^{n-1}\times(0,+\infty)\to\mathbb{R}^{n}\setminus\{0\}$) given by $\rho_{\infty}^{-1}(x)=(\frac{x}{\|x\|},\frac{1}{\|x\|})$ (resp. $\rho_{p}^{-1}(x)=(\frac{x-p}{\|x-p\|},\|x-p\|)$).
The strict transform of the subset $X$ under the spherical blowing-up $\rho_{\infty}$ is $X^{\prime}_{\infty}:=\overline{\rho_{\infty}^{-1}(X\setminus\{0\})}$ (resp. $X^{\prime}_{p}:=\overline{\rho_{p}^{-1}(X\setminus\{0\})}$). The subset $X_{\infty}^{\prime}\cap(\mathbb{S}^{n-1}\times\{0\})$ (resp. $X_{p}^{\prime}\cap(\mathbb{S}^{n-1}\times\{0\})$) is called the boundary of $X^{\prime}_{\infty}$ (resp. $X^{\prime}_{p}$) and it is denoted by $\partial X^{\prime}_{\infty}$ (resp. $\partial X^{\prime}_{p}$).
Remark 2.11.
If $X\subset\mathbb{R}^{n}$ is a definable set in an o-minimal structure, then $\partial X^{\prime}_{\infty}=(C(X,\infty)\cap\mathbb{S}^{n-1})\times\{0\}$ (resp. $\partial X^{\prime}_{p}=(C(X,p)\cap\mathbb{S}^{n-1})\times\{0\}$.
We finish this Subsection by reminding the invariance of the tangent cone at infinity under bi-Lipschitz homeomorphisms at infinity obtained in the paper [39] (see also [14, Theorem 2.19]).
Theorem 2.12 (Corollary 2.11 in [39]).
Let $X\subset\mathbb{R}^{n}$ and $Y\subset\mathbb{R}^{m}$ be unbounded definable sets with $d=\dim X=\dim Y$. If there exist compact subsets $K\subset\mathbb{R}^{n}$ and $\tilde{K}\subset\mathbb{R}^{m}$, and a bi-Lipschitz homeomorphism $\varphi\colon X\setminus K\to Y\setminus\tilde{K}$, then there is a globally defined bi-Lipschitz homeomorphism $d\varphi\colon C(X,\infty)\to C(Y,\infty)$ with $d\varphi(0)=0$.
2.6. Relative multiplicities at infinity
Let $X\subset\mathbb{R}^{n}$ be a $d$-dimensional definable subset in an o-minimal structure and $p\in\mathbb{R}^{n}\cup\{\infty\}$. We say $x\in\partial X^{\prime}_{p}$ is simple point of $\partial X^{\prime}_{p}$, if there is an open subset $U\subset\mathbb{R}^{n+1}$ with $x\in U$ such that:
  • a)
    the connected components $X_{1},\cdots,X_{r}$ of $(X^{\prime}_{p}\cap U)\setminus\partial X^{\prime}_{p}$ are topological submanifolds of $\mathbb{R}^{n+1}$ with $\dim X_{i}=\dim X$, for all $i=1,\cdots,r$;
  • b)
    $(X_{i}\cup\partial X^{\prime}_{p})\cap U$ are topological manifolds with boundary $\partial X^{\prime}_{p}\cap U$, for all $i=1,\cdots,r$.
Let ${\rm Smp}(\partial X^{\prime}_{p})$ be the set of simple points of $\partial X^{\prime}_{p}$ and we define $C_{{\rm Smp}}(X,p)=\{t\cdot x;\,t>0\mbox{ and }(x,0)\in{\rm Smp}(\partial X^{\prime})\}$. Let $k_{X}\colon C_{{\rm Smp}}(X,p)\to\mathbb{N}$ be the function such that $k_{X,p}(x)$ is the number of connected components of the germ $(\rho^{-1}(X\setminus\{p\}),v)$, where $v=(\frac{x}{\|x\|},0)$.
Remark 2.13.
${\rm Smp}(\partial X^{\prime}_{p})$ is an open dense subset of the $(d-1)$-dimensional part of $\partial X^{\prime}_{p}$ whenever $\partial X^{\prime}_{p}$ is a $(d-1)$-dimensional subset, where $d=\dim X$.
Definition 2.14.
It is clear that the function $k_{X,p}$ is locally constant. In fact, $k_{X,p}$ is constant on each connected component $X_{j}$ of $C_{{\rm Smp}}(X,p)$. Then, we define the relative multiplicity at $p$ of $X$ (along of $X_{j}$) to be $k_{X,p}(X_{j}):=k_{X,p}(x)$ with $x\in X_{j}$.
Definition 2.15.
Let $X\subset\mathbb{R}^{n}$ and $Y\subset\mathbb{R}^{m}$ be closed sets. Let $p\in\mathbb{R}^{n}\cup\{\infty\}$, $q\in\mathbb{R}^{m}\cup\{\infty\}$ and a homeomorphism $\varphi:X\rightarrow Y$ such that $q=\lim\limits_{x\rightarrow p}{\varphi(x)}$, is said a blow-spherical homeomorphism at $p$, if the homeomorphism
$\varphi^{\prime}\colon\rho_{p}^{-1}(X\setminus\{p\})\rightarrow\rho_{q}^{-1}(Y\setminus\{q\})$
extends to a homeomorphism $\varphi^{\prime}\colon X^{\prime}_{p}\rightarrow Y^{\prime}_{q}$.
2.7. Blow-spherical invariance of the relative multiplicities at infinity
The following result was proved in [40, Proposition 3.5] for semialgebraic sets, but the proof is the same for definable sets in an o-minimal structure.
Proposition 2.16.
Let $X\subset\mathbb{R}^{n}$ and $Y\subset\mathbb{R}^{m}$ be definable subsets in an o-minimal structure on $\mathbb{R}$. Let $\varphi\colon X\rightarrow Y$ be a blow-spherical homeomorphism at $p\in\mathbb{R}^{n}\cup\{\infty\}$. Then $\varphi^{\prime}(Smp(\partial X^{\prime}_{p}))=Smp(\partial Y^{\prime}_{q})$ and
$k_{X,p}(x)=k_{Y,q}(\varphi^{\prime}(x)),$
for all $x\in Smp(\partial X^{\prime}_{p})$, where $q=\lim\limits_{x\rightarrow p}{\varphi(x)}$.
2.8. Bi-Lipschitz invariance of the relative multiplicities at infinity
Theorem 2.17.
Let $X\subset\mathbb{R}^{n}$ and $Y\subset\mathbb{R}^{m}$ be unbounded definable sets with $d=\dim X=\dim Y$. If there exist compact subsets $K\subset\mathbb{R}^{n}$ and $\tilde{K}\subset\mathbb{R}^{m}$, and a bi-Lipschitz homeomorphism $\varphi\colon X\setminus K\to Y\setminus\tilde{K}$, then there exists a bi-Lipschitz homeomorphism $d\varphi\colon C(X,\infty)\to C(Y,\infty)$ that satisfies
$k_{X,\infty}(x)=k_{Y,\infty}(d\varphi(x)),\quad\forall x\in C_{{\rm Smp}}(X,\infty)\cap d\varphi^{-1}(C_{{\rm Smp}}(Y,\infty)).$
In particular, $d\varphi|_{C_{{\rm Smp}}(X,\infty)}\colon C_{{\rm Smp}}(X,\infty)\to C_{{\rm Smp}}(Y,\infty)$ is a bi-Lipschitz homeomorphism, whenever $C_{{\rm Smp}}(X,\infty)\not=\emptyset$.
Proof.
We follow closely the proof of Theorem 4.2 in [37], but because its importance in this article, we present the proof here.
By making the identifications $X\setminus K\longleftrightarrow X\setminus K\times\{0\}$ and $Y\setminus\tilde{K}\longleftrightarrow\{0\}\times Y\setminus\tilde{K}$, we can assume that $m=n$ and $\varphi$ is globally defined in $\mathbb{R}^{n}$. We know that there exist a sequence of positive real numbers $S=\{t_{k}\}_{k\in\mathbb{N}}$ and a bi-Lipschitz homeomorphism $d\varphi\colon C(X,\infty)\to C(Y,\infty)$ such that
$t_{k}\to+\infty\quad\mbox{and}\quad\frac{\varphi(t_{k}v)}{t_{k}}\to d\varphi(v),\quad\forall v\in C(X,\infty).$
For more details, see [35], [14], [15] and [39].
Thus, $\rho^{-1}\circ\varphi\circ\rho\colon SX\rightarrow Y^{\prime}$ is an injective and continuous mapping that extends continuously to a mapping $\varphi^{\prime}\colon\overline{SX}\to Y^{\prime},$ where $\rho=\rho_{\infty}$ and
$SX=\{(x,s)\in\mathbb{S}^{n-1}\times(0,+\infty);\,\textstyle\frac{1}{s}\cdot x\in X\mbox{ and }\frac{1}{s}\in S\}.$
We remark that $Smp(\partial X^{\prime}_{\infty})=\emptyset$ (resp. $Smp(\partial Y^{\prime}_{\infty})=\emptyset$) if and only if $\dim C(X,\infty)<\dim X$ (resp. $\dim C(Y,\infty)<\dim Y$). Since by Theorem 2.12, $\dim C(X,{\infty})=\dim C(Y,{\infty})$, then we obtain that $Smp(\partial X^{\prime}_{\infty})=\emptyset$ if and only if $Smp(\partial Y^{\prime}_{\infty})=\emptyset$. However, when $Smp(\partial X^{\prime}_{\infty})=Smp(\partial Y^{\prime}_{\infty})=\emptyset$, we have nothing to do. Thus, we can assume that $Smp(\partial X^{\prime}_{\infty})\not=\emptyset$ and, thus, $Smp(\partial Y^{\prime}_{\infty})\not=\emptyset$ as well. In particular, $C_{{\rm Smp}}(X,{\infty})$ (resp. $C_{{\rm Smp}}(Y,{\infty})$) is a dense subset in the $d$-dimensional part of $C(X,{\infty})$ (resp. $C(Y,{\infty})$). Then, $d\varphi(C_{{\rm Smp}}(X,{\infty}))\cap C_{{\rm Smp}}(Y,{\infty})$ is a dense subset in the $d$-dimensional part of $C(Y,{\infty})$ and a dense subset in $C_{{\rm Smp}}(Y,{\infty})$.
Let $X_{1},\cdots,X_{r}$ be the connected components of $C_{{\rm Smp}}(X,{\infty})$ and let $Y_{1},\cdots,Y_{s}$ be the connected components of $C_{{\rm Smp}}(Y,{\infty})$. For each point $x\in X_{j}$, we know that $k_{X,{\infty}}(X_{j})=k_{X,{\infty}}(x)$ is the number of connected components of the germ $(\rho^{-1}_{\infty}(X)\cap B_{\delta}(x),x)$. Then, $k_{X,{\infty}}(x)$ can be seen as the number of connected components of the set $(SX\cap\mathbb{S}^{n}\times\{t_{k}\})\cap B_{\delta}(x)$ for all large enough $k$.
Let $\pi\colon\mathbb{R}^{n}\to\mathbb{R}^{d}$ be a linear projection such that
$\pi^{-1}(0)\cap(C(X,{\infty})\cup C(Y,{\infty}))=\{0\}.$
Let us denote the ramification locus of
$\pi_{|X}\colon X\to\mathbb{R}^{d}\quad\mbox{and}\quad\pi_{|C(X,{\infty})}\colon C(X,{\infty})\to\mathbb{R}^{d}$
by $\sigma(X)$ and $\sigma(C(X,{\infty}))$ respectively. Similarly, we define $\sigma(Y)$ and $\sigma(C(Y,{\infty}))$.
Let $Z=\sigma(X)\cup\sigma(C(X,{\infty}))$, $W=\sigma(Y)\cup\sigma(C(Y,{\infty}))$ and
$\Sigma=Z\cup C(Z,{\infty})\cup\pi(d\varphi^{-1}(\pi|_{Y\cup C(Y,{\infty})}^{-1}(W\cup C(W,{\infty}))).$
Since $\pi|_{Y\cup C(Y,{\infty})}$ is a proper mapping, $d\varphi$ is a bi-Lipschitz homeomorphism and $\dim Z\cup W\leq d-1$, we obtain that $\dim\Sigma\leq d-1$. Thus, we obtain that if $v^{\prime}\in\mathbb{R}^{d}\setminus\Sigma$ then $\pi|_{X}^{-1}(v^{\prime})\subset\mathbb{R}^{n}\setminus\pi|_{X}^{-1}(Z\cup C(Z,{\infty}))$ and for any $v_{i}\in\pi|_{X}^{-1}(v^{\prime})$, we have that $w^{\prime}_{i}=\pi(d\varphi(v_{i}))\in\mathbb{R}^{p}\setminus(W\cup C(W,{\infty}))$. For $\eta,R>0$, we define the following set
$C_{\eta,R}(v^{\prime})=\{w\in\mathbb{R}^{d}|\,\exists t>0;\|tv^{\prime}-w\|<\eta t\}\setminus B_{R}(0).$
Let $\eta,R>0$ be sufficiently big such that $C_{\eta,R}(v^{\prime})\subset\mathbb{R}^{d}\setminus(Z\cup C(Z,{\infty}))$. Then, we obtain that $\pi|_{V}\colon V\to C_{\eta,R}(v^{\prime})$ is a bi-Lipschitz homeomorphism for each connected component $V$ of $\pi^{-1}(C_{\eta,R}(v^{\prime}))\cap X$. Therefore, for each $j=1,\dots,r$, there are different connected components $V_{j1},\dots,V_{jk_{X}(X_{j})}$ of $\pi^{-1}(C_{\eta,R}(v^{\prime}))\cap X$ such that
$C(\overline{V_{ji}},{\infty})\setminus\{0\}\subset X_{j},\quad\forall i\in\{1,...,k_{X,{\infty}}(X_{j})\}.$
Let $\rho_{\infty}\colon\mathbb{S}^{n-1}\times(0,+\infty)\to\mathbb{R}^{n}$ be the spherical blowing-up at infinity. Fixed $j\in\{1,\cdots,r\}$, let us suppose that there is $q\in\{1,\cdots,s\}$ such that $d\varphi(X_{j})\cap Y_{q}\not=\emptyset$ and $k_{X,\infty}(X_{j})>k_{Y,\infty}(Y_{q})$, it means that if we consider a point $x=(v,0)\in(X_{j}\cap\mathbb{S}^{n-1})\times\{0\}$ with $d\varphi(x)\in Y_{q}$ and, since $\dim\Sigma\leq d-1$, we can assume that $\pi(v)\not\in\Sigma$, then there are at least two different connected components $V_{ji}$ and $V_{jl}$ of $\pi^{-1}(C_{\eta,R}(\pi(v)))\cap X$ and sequences $\{(x_{k},t_{k}^{-1})\}_{k\in\mathbb{N}}\subset\rho^{-1}(V_{ji})\cap SX$ and $\{(y_{k},t_{k}^{-1})\}_{k\in\mathbb{N}}\subset\rho^{-1}(V_{jl})\cap SX$ such that $\lim(x_{k},t_{k}^{-1})=\lim(y_{k},t_{k}^{-1})=x$ and $\varphi^{\prime}(x_{k},t_{k}^{-1}),\varphi^{\prime}(y_{k},t_{k}^{-1})\in\rho^{-1}(\widetilde{V}_{qm})$ for all $k\in\mathbb{N}$, where $\widetilde{V}_{qm}$ is a connected component of $\pi^{-1}(C_{\tilde{\eta},\tilde{R}}(\pi(d\varphi(v))))\cap Y$ for some $\tilde{\eta},\tilde{R}>0$ such that
$C_{\tilde{\eta},\tilde{R}}(w^{\prime})=\{\tilde{w}\in\mathbb{R}^{p}|\,\exists t>0;\|tw^{\prime}-\tilde{w}\|<\tilde{\eta}t\}\setminus B_{\tilde{R}}(0)$
satisfies $C_{\tilde{\eta},\tilde{R}}(w^{\prime})\subset\mathbb{R}^{d}\setminus(W\cup C(W,{\infty}))$, where $w^{\prime}=\pi(d\varphi(v))$. In particular, $\pi|_{\widetilde{V}}\colon\widetilde{V}\to C_{\tilde{\eta},\tilde{R}}(w^{\prime})$ is a bi-Lipschitz homeomorphism, where $\widetilde{V}=\widetilde{V}_{qm}$. Since $\varphi(t_{k}x_{k}),\varphi(t_{k}y_{k})\in\widetilde{V}$ for all $k\in\mathbb{N}$, we have
$\|\varphi(t_{k}x_{k})-\varphi(t_{k}y_{k})\|=o(t_{k})$
and
$d_{Y,in}(\varphi(t_{k}x_{k}),\varphi(t_{k}y_{k}))\leq d_{\widetilde{V},in}(\varphi(t_{k}x_{k}),\varphi(t_{k}y_{k}))=o(t_{k}).$
Now, since $X$ is bi-Lipschitz homeomorphic at infinity to $Y$, we obtain $d_{X,in}(t_{k}x_{k},t_{k}y_{k})\leq o_{\infty}(t_{k})$. On the other hand, since $t_{k}x_{k}$ and $t_{k}y_{k}$ lie in different connected components of $\pi^{-1}(C_{\eta,R}(\pi(v)))\cap X$, there exists a constant $C>0$ such that $d_{X,in}(t_{k}x_{k},t_{k}y_{k})\geq Ct_{k}$, which is a contradiction.
We have proved that $k_{X,{\infty}}(X_{j})\leq k_{Y,{\infty}}(Y_{q})$. By similar arguments, by using that $\varphi^{-1}$ is a bi-Lipschitz homeomorphism, we can also prove $k_{Y,{\infty}}(Y_{q})\leq k_{X,{\infty}}(X_{j})$. Therefore, $k_{X,{\infty}}(X_{j})=k_{Y,{\infty}}(Y_{q})$ for any $q\in\{1,\cdots,s\}$ such that $d\varphi(Cone^{*}(X_{j}))\cap Cone^{*}(Y_{q})\not=\emptyset$. ∎
2.9. Density at infinity
Let $X\subset\mathbb{R}^{m}$ be a definable set. Fixed $n\in\mathbb{N}$ and $p\in X$, we define the function $\theta^{n}(X,p,\cdot)\colon(0,+\infty)\to X$ by the following:
$\theta^{n}(X,p,r)=\frac{\mathcal{H}^{n}(X\cap B_{r}^{n}(p))}{\mu_{n}r^{n}},$
where $\mu_{n}$ is the volume of the $n$-dimensional Euclidean unit ball, $\mathcal{H}^{n}(A)$ is the $n$-dimensional Hausdorff measure of $A$ and $B_{r}^{m}(p)\subset\mathbb{R}^{m}$ is the open Euclidean ball center at $p$ of radius $r>0$.
By simplicity, we denote $B_{r}^{m}:=B_{r}^{m}(0)$.
Definition 2.18.
Let $X\subset\mathbb{R}^{n+k}$ be set. We say that $X$ has $n$-dimensional density at $p$ and we denote it by $\theta(X,p)$, if the limit exists:
$\theta^{n}(X,p):=\lim_{r\rightarrow 0^{+}}{\theta(X,p,r)}.$
Definition 2.19.
Let $X\subset\mathbb{R}^{n+k}$ be a set. We say that $X$ has $n$-dimensional density at infinity and we denote it by $\theta^{n}(X)$, if the limit exists:
$\theta^{n}(X):=\lim_{r\rightarrow+\infty}\frac{\mathcal{H}^{n}(X\cap B_{r}^{n+k})}{\mu_{n}r^{n}}.$
When $X$ has $n$-dimensional density at infinity and $n=\dim X$, we just say that $X$ has density at infinity.
3. Kurdyka-Raby’s formula at infinity
The goal of this section is to bring the Theorem of Kurdyka and Raby in [23] and its ideas to a global perspective. Then, we prove the following Kurdyka-Raby’s formula at infinity:
Theorem 3.1 (Kurdyka-Raby’s formula at infinity).
Let $X\subset\mathbb{\mathbb{R}}^{n+k}$ be a definable set in an o-minimal structure on $\mathbb{R}$ and $n=\dim X$. Let $C_{1},...,C_{m}$ be the connected components of $C_{{\rm Smp}}(X,\infty)$. Then, $\theta^{n}(X)$ is defined and for each $x\in\mathbb{R}^{n+k}$, we have
$\theta^{n}(X)=\lim\limits_{r\to+\infty}\frac{\mathcal{H}^{n}(X\cap B_{r}^{n+k}(x))}{\mu_{n}r^{n}}=\sum\limits_{j=1}^{m}k_{X,\infty}(C_{j})\frac{\mathcal{H}^{n}(C_{j}\cap B_{1}^{n+k})}{\mu_{n}}.$
3.1. Proof of the Kurdyka-Raby’s formula at infinity
Proof of Theorem 3.1.
We first prove that the result is true for open definable sets.
Claim 3.1.1.
Let $\Omega\subset\mathbb{R}^{n}$ be an open definable set. Then, the density at infinity satisfies:
$\lim_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}\left(\Omega\cap B_{r}^{n}\right)}{r^{n}}}=\mathcal{H}^{n}\left(B_{1}^{n}\cap C(\Omega,\infty)\right).$
Proof of the Claim 3.1.1.
First, note that $\theta^{n}(\Omega)$ there exists if and only if there exists $\theta^{n}(\Omega\setminus K)$, where $K\subset\Omega$ is a compact subset. Moreover, in this case, $\theta^{n}(\Omega)=\theta^{n}(\Omega\setminus K)$ In fact, for every $R>0$, we have:
$\theta^{n}(\Omega)$ $=$ $\lim_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}(\Omega\cap B_{r}^{n})}{\mu_{n}r^{n}}}$
$=$ $\lim_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}(\Omega\cap K\cap B^{n}_{r})+\mathcal{H}^{n}((\Omega\setminus K)\cap B^{n}_{r})}{\mu_{n}r^{n}}}$
$=$ $\lim_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}((\Omega\setminus K)\cap B^{n}_{r})}{\mu_{n}r^{n}}}$
$=$ $\theta^{n}(\Omega\setminus K).$
When $C$ is the cone with the vertex at 00 in $\mathbb{R}^{n}$, we have
$\mathcal{H}^{n}\left(C\cap B_{r}^{n}\right)=r^{n}\mathcal{H}^{n}\left(C\cap B^{n}_{1}\right),$
for every $r>0$. Now, we consider the cone $C_{r}:=\mbox{Cone}_{0}\left(\Omega\setminus B_{r}^{n}\right)$.
Claim 3.1.2.
$\bigcap_{r>0}{\overline{C_{r}}}=C(\Omega,\infty)$.
Proof of the Claim 3.1.2.
Let $v\in\bigcap_{r>0}{\overline{C_{r}}}$. Then, $v\in C_{k}$ for every $l\in\mathbb{N}$. Therefore, $v=t_{l}x_{l}$, where $x_{l}\in\Omega\setminus B^{n}_{l}$, with $\|x_{l}\|\geq l$, and $t_{l}\in(0,+\infty)$. Hence, $v\in C(\Omega,\infty)$.
Conversely, if $v\in C(\Omega,\infty)$, it follows from [14, Proposition 2.15] that there exists a continuous and definable curve $\alpha\colon(\epsilon,+\infty)\rightarrow\Omega$ such that $\alpha(t)=tv+o(t)$ with $\|\alpha(t)\|>t$. Now, define $\beta(t)=v+\frac{o(t)}{t}$, and note that $\|\beta(t)-v\|\rightarrow 0$ as $t\rightarrow+\infty$. Since $\alpha(t)\subset C_{t}$, we have $\frac{\alpha(t)}{t}\in C_{t}$, and therefore, $\beta(t)\in C_{t}$. Thus,
$\text{dist}(v,C_{t})\leq\|v-\beta(t)\|,$
for every $t\geq 0$. However, $C_{t}\subset C_{s}$ whenever $s<t$. Then for any $s>0$ and $s<t$, we have
$\text{dist}(v,\overline{C_{s}})\leq\text{dist}(v,\overline{C_{t}})\leq\|v-\beta(t)\|\rightarrow 0,$
as $t\rightarrow+\infty$. Hence, $v\in\overline{C_{s}}$ for any $s>0$ and, consequently, $v\in\bigcap_{r>0}{\overline{C_{r}}}$. ∎
Now, consider $Z=\bigcap_{r>0}{C_{r}}$, and it follows from the claim that $C(\Omega,\infty)\supset Z$.
Moreover, the functions $r\mapsto\mathcal{H}^{n}\left(C_{r}\cap B_{1}^{n}\right)$ and $r\mapsto\mathcal{H}^{n}\left(\overline{C_{r}}\cap B_{1}^{n}\right)$ are non-increasing; hence, the following limits exist:
$\lim_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}\left(C_{r}\cap B^{n}_{r}\right)}{r^{n}}}$ $=$ $\lim_{r\rightarrow+\infty}{\mathcal{H}^{n}\left(C_{r}\cap B^{n}_{1}\right)}$
$=$ $\mathcal{H}^{n}\left(B^{n}_{1}\cap\bigcap_{r>0}C_{r}\right)$
$=$ $\mathcal{H}^{n}\left(B^{n}_{1}\cap Z\right)$
and
$\lim_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}\left(\overline{C_{r}}\cap B^{n}_{r}\right)}{r^{n}}}$ $=$ $\lim_{r\rightarrow+\infty}{\mathcal{H}^{n}\left(\overline{C_{r}}\cap B^{n}_{1}\right)}$
$=$ $\mathcal{H}^{n}\left(B^{n}_{1}\cap\bigcap_{r>0}\overline{C_{r}}\right)$
$=$ $\mathcal{H}^{n}(B^{n}_{1}\cap C(\Omega,\infty)).$
Since $C_{r}$ is a definable set, we have
$\mathcal{H}^{n}\left(B_{1}^{n}\cap C_{r}\right)=\mathcal{H}^{n}\left(B_{1}^{n}\cap\overline{C}_{r}\right).$
Therefore, $\mathcal{H}^{n}(B^{n}_{1}\cap C(\Omega,\infty))=\mathcal{H}^{n}(B^{n}_{1}\cap Z)$.
Moreover, for every $R>0$, we have $C_{R}\supset\Omega\setminus(\Omega\cap B^{n}_{R})$. Thus,
$\limsup_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}\left(\Omega\cap B^{n}_{r}\right)}{r^{n}}}$ $=$ $\limsup_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}\left((\Omega\setminus(\Omega\cap B^{n}_{R}))\cap B^{n}_{r}\right)}{r^{n}}}$
$\leq$ $\lim_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}\left(C_{r}\cap B^{n}_{r}\right)}{r^{n}}}$
$=$ $\mathcal{H}^{n}(Z\cap B^{n}_{1})=\mathcal{H}^{n}(B^{n}_{1}\cap C(\Omega,\infty)).$
For $w\in\mathbb{R}^{n}\setminus\{0\}$, we set $(w,+\infty):=\{tw;t>1\}$. Now, take $z\in Z\setminus\{0\}$. We assume without loss of generality that $0\not\in\Omega$. Then, $(z,+\infty)\cap\Omega$ is a definable set in $\Omega$, and we define $\alpha(z)=\sup\left\{\lambda\geq 0;\lambda z\notin\Omega\right\}$, which is non-negative. If $\alpha(z)\not=0$, the segment $(\alpha(z)z,+\infty)$ is contained in $\Omega$ and if $\alpha(z)=0$, then $(\epsilon z,+\infty)$ is contained in $\Omega$, for all $\epsilon>0$.
We define the set
$Z_{r}=\{z\in Z\setminus\{0\}/\|\alpha(z)\cdot z\|\leq r\}.$
Since $\alpha(\mu z)=\sup\{\lambda\geq 0;\lambda(\mu z)\notin\Omega\}=\lambda_{0}$ and $\alpha(z)=\sup\{\eta\geq 0;\eta z\notin\Omega\}=\eta_{0}$, we have $\lambda_{0}=\frac{\eta_{0}}{\mu}$, and $Z_{r}$ is indeed a cone.
Now, note that $Z_{r}\setminus\overline{B^{n}_{r}}\subset\Omega\setminus\overline{B^{n}_{r}}$. Indeed, if $z\in Z_{r}\setminus\overline{B^{n}_{r}}$, then $\|z\|>r$, and additionally, $\|\alpha(z)z\|\leq r$. Therefore, $\alpha(z)\leq\frac{r}{\|z\|}<1$. Note that for $r<s$, we have $Z_{r}\subset Z_{s}$. In fact, if $z\in Z_{r}$, then $\alpha(z)\|z\|\leq r<s$, so $z\in Z_{s}$.
Claim 3.1.3.
$Z\setminus\{0\}=\bigcup_{r>0}{Z_{r}}$.
Proof.
It follows from the definition that $\bigcup_{r>0}{Z_{r}}\subset Z\setminus\{0\}$. Reciprocally, assume that $z\in Z\setminus\{0\}$. Thus for each $r>0$ there exists a points $z_{r}\in\Omega\setminus B^{n}_{r}$ and $\lambda_{r}\in(0,+\infty)$ such that $\lambda_{r}z_{r}=z$ and since $\Omega\setminus\overline{B^{n}_{r}}$ is open set there exists $t_{r_{j}}\in(0,+\infty)$ such that $t_{r_{j}}z_{j}\in\Omega\setminus\overline{B^{n}_{r}}$. So, if the semi-line $\vec{0z}$ we have $(t_{r_{j}}z_{j},+\infty):=\vec{0z}\setminus(\vec{0z}\cap B^{n}_{t_{r_{j}}})\subset\Omega$, we have $\alpha(z)<+\infty$. Finally, note that $(z,+\infty)\subset\Omega\setminus B_{r}^{n}$.
Moreover, for $r>R>0$, we have
$\liminf_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}(\Omega\cap B^{n}_{r})}{r^{n}}}$ $=$ $\liminf_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}((\Omega\setminus B^{n}_{R})\cap B^{n}_{r})}{r^{n}}}$
$\geq$ $\liminf_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}((Z_{r}\setminus B^{n}_{R})\cap B^{n}_{r})}{r^{n}}}$
$=$ $\liminf_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}(Z_{r}\cap B^{n}_{r})-\mathcal{H}^{n}(Z_{r}\cap B^{n}_{R})}{r^{n}}}$
$=$ $\liminf_{r\rightarrow+\infty}{\left(\mathcal{H}^{n}(Z_{r}\cap B^{n}_{1})-\frac{\mathcal{H}^{n}(Z\cap B^{n}_{R})}{r^{n}}\right)}$
$=$ $\mathcal{H}^{n}(Z\cap B^{n}_{1})=\mathcal{H}^{n}(B^{n}_{1}\cap C(\Omega,\infty)).$
Therefore, $\Omega$ has density at infinity and
$\theta^{n}(\Omega)=\frac{\mathcal{H}^{n}(B^{n}_{1}\cap C(\Omega,\infty))}{\mu_{n}}=\theta^{n}(C(\Omega,\infty)).$
Now we prove that $X$ has density at infinity. Recently, this was proved in [29], but here we present a different proof.
Claim 3.1.4.
$X$ has density at infinity.
Proof of the Claim 3.1.4.
Since the tangent mapping $\nu\colon\mathcal{M}:=X\setminus{\rm Sing}_{1}(X)\to G_{n}(\mathbb{R}^{n+k})$, given by $\nu(x)=T_{x}\mathcal{M}$, is a definable mapping (see [22, Lemma 1.14]), likewise it was done in [44], we obtain that for each $\epsilon>0$, there exist $N(\epsilon)$ and disjoint definable sets $\Gamma_{1}^{\epsilon},\ldots,\Gamma_{N(\epsilon)}^{\epsilon}$ contained in $X$ such that:
  1. (1)
    $\dim X\setminus\cup_{i=1}^{N(\epsilon)}\Gamma_{i}^{\epsilon}<n$;
  2. (2)
    Each $\Gamma_{i}^{\epsilon}$ is the graph of a mapping $\varphi_{i}^{\epsilon}\colon U_{i}^{\epsilon}\rightarrow(E_{i}^{\epsilon})^{\bot}$ with derivative bounded by $\epsilon$ and $U_{i}^{\epsilon}$ is a definable open set in $E_{i}^{\epsilon}\in G_{n}(\mathbb{R}^{n+k})$, where $G_{n}(\mathbb{R}^{n+k})$ is the Grassmannian of all $n$-dimensional linear subspaces of $\mathbb{R}^{n+k}$.
By refining the above decomposition, if necessary, we may assume that each $U_{i}^{\epsilon}$ is $M$-LNE, where $M=M(n)$ depends only on $n$ (see, e.g., [22, Theorem 1.2]). Therefore, each $\varphi_{i}^{\epsilon}$ is $M\epsilon$-Lipschitz. Thus, the mapping $\psi_{i}^{\epsilon}\colon U_{i}^{\epsilon}\to\Gamma_{i}^{\epsilon}$, given by $\psi_{i}^{\epsilon}(x)=x+\varphi_{i}^{\epsilon}(x)$ is a definable lipeomorphism such that
$\|x-y\|\leq\|\psi_{i}^{\epsilon}(x)-\psi_{i}^{\epsilon}(y)\|\leq(1+M\epsilon)\|x-y\|\quad\forall x,y\in U_{i}^{\epsilon}.$
By Claim 3.1.1, it follows that each $U_{i}^{\epsilon}$ has a density at infinity and
$\theta^{n}(U_{i}^{\epsilon})=\lim\limits_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}(U_{i}^{\epsilon}\cap B_{r}^{n})}{\mu_{n}r^{n}}}=\frac{\mathcal{H}^{n}(B_{1}^{n}\cap C(U_{i}^{\epsilon},\infty))}{\mu_{n}}=\theta^{n}(C(U_{i}^{\epsilon},\infty)).$
Thus,
$\frac{\theta^{n}(U_{i}^{\epsilon})}{(1+M\epsilon)^{n}}\leq\liminf_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}(\Gamma_{i}^{\epsilon}\cap B_{r}^{n})}{\mu_{n}r^{n}}}\leq\limsup_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}(\Gamma_{i}^{\epsilon}\cap B_{r}^{n})}{\mu_{n}r^{n}}}\leq(1+M\epsilon)^{n}\theta^{n}(U_{i}^{\epsilon}).$
By setting $\lambda(\epsilon)=\sum_{i=1}^{N(\epsilon)}\theta^{n}(U_{i}^{\epsilon})$ and since for every $r>0$, we have
$\mathcal{H}^{n}(X\cap B_{r}^{n+k})=\sum_{i=1}^{N(\epsilon)}{\mathcal{H}^{n}(\Gamma_{i}^{\epsilon}\cap B_{r}^{n+k})},$
then we obtain
$\frac{\lambda(\epsilon)}{(1+M\epsilon)^{n}}\leq\liminf_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}(X\cap B_{r}^{n})}{\mu_{n}r^{n}}}\leq\limsup_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}(X\cap B_{r}^{n})}{\mu_{n}r^{n}}}\leq(1+M\epsilon^{\prime})^{n}\lambda(\epsilon^{\prime}),$
for all $\epsilon,\epsilon^{\prime}>0$.
Note that $\lambda(\epsilon)\leq(1+M\epsilon)^{n}2^{n}\lambda(1)$. Then
$\lim_{\epsilon\rightarrow 0}{(1+M\epsilon)^{n}\lambda(\epsilon)-\frac{\lambda(\epsilon)}{(1+M\epsilon)^{n}}}=0.$
Therefore $\liminf_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}(X\cap B_{r}^{n})}{\mu_{n}r^{n}}}=\limsup_{r\rightarrow+\infty}{\frac{\mathcal{H}^{n}(X\cap B_{r}^{n})}{\mu_{n}r^{n}}}$, and thus $\theta^{n}(X)$ is defined.
In this case, it is easy to see that
$\theta^{n}(X)=\lim\limits_{r\to+\infty}\frac{\mathcal{H}^{n}(X\cap B_{r}^{n+k}(x))}{\mu_{n}r^{n}}$
for all $x\in\mathbb{R}^{n+k}$.
Finally, we are going to prove the second part, more precisely,
$\theta^{n}(X)=\sum\limits_{j=1}^{m}k_{X,\infty}(C_{j})\cdot\mathcal{H}^{n}(C_{j}),$
where $C_{1},...,C_{m}$ are the connected components of $C_{{\rm Smp}}(X,\infty)$.
It is worth noting that the above formula is not presented in [29].
Fixed $\epsilon>0$ and $i\in\{1,...,N(\epsilon)\}$, let $U=U_{i}^{\epsilon}$, $\Gamma=\Gamma_{i}^{\epsilon}$, $\varphi=\varphi_{i}^{\epsilon}$ and $E=E_{i}^{\epsilon}$. Therefore, it follows from Theorems 2.12 and 2.17 that
  1. (1)
    $C(\Gamma,\infty)$ is the graph of a definable $M\epsilon$-Lipschitz mapping $d_{\infty}\varphi\colon C(U,\infty)\to E^{\perp}$;
  2. (2)
    $k_{\Gamma,\infty}(x)=1$ for all $x\in C_{\rm Smp}(\Gamma,\infty)$.
Then
$\frac{1}{(1+M\epsilon)^{n}}\theta^{n}(U_{i}^{\epsilon})\leq\theta^{n}\left(\Gamma_{i}^{\epsilon}\right)\leq(1+M\epsilon)^{n}\theta^{n}(U_{i}^{\epsilon}).$
and
$\frac{1}{(1+M\epsilon)^{n}}\theta^{n}(C(U_{i}^{\epsilon},\infty))\leq\theta^{n}\left(C(\Gamma_{i}^{\epsilon},\infty)\right)\leq(1+M\epsilon)^{n}\theta^{n}(C(U_{i}^{\epsilon},\infty)).$
Again, according to Claim 3.1.1, we have
$\theta^{n}(C_{\rm Smp}(U_{i}^{\epsilon},\infty))=\theta^{n}(C(U_{i}^{\epsilon},\infty))=\theta^{n}\left(U_{i}^{\epsilon}\right).$
Therefore,
(1) $\frac{1}{(1+M\epsilon)^{2n}}\theta^{n}(C(\Gamma_{i}^{\epsilon},\infty))\leq\theta^{n}\left(\Gamma_{i}^{\epsilon}\right)\leq(1+M\epsilon)^{2n}\theta^{n}(C(\Gamma_{i}^{\epsilon},\infty)).$
So, let $\mathcal{A}$ be a stratification of $C_{\rm Smp}(X,\infty)$ compatible with $\mathbb{S}^{{n+k}-1}$, $C_{1},\cdots,C_{m},C_{\rm Smp}(\Gamma_{1}^{\epsilon},\infty),\cdots,C_{\rm Smp}(\Gamma_{N(\epsilon)}^{\epsilon},\infty)$.
For all $1\leq i\leq N(\epsilon)$, $k_{\Gamma_{i}^{\epsilon},\infty}(x)=1$ for all $x\in C_{\rm Smp}(\Gamma_{i}^{\epsilon},\infty)$. Therefore, if $T\in\mathcal{A}$ is a stratum of dimension $n-1$ contained in $C_{j}\cap\mathbb{S}^{{n+k}-1}$, it follows from the definition of $k_{X,\infty}(C_{j})$ that it is the number of pieces $\Gamma_{i}^{\epsilon}$ such that $T$ is contained in $C_{\rm Smp}(\Gamma_{i}^{\epsilon})$. Then
$\sum_{i=1}^{N(\epsilon)}{\mathcal{H}^{n}(C_{\rm Smp}(\Gamma_{i}^{\epsilon})\cap B^{{n+k}})}=\sum_{j=1}^{m}{k_{X,\infty}(C_{j})}\mathcal{H}^{n-1}(C_{j}\cap B^{{n+k}}).$
and thus
(2) $\sum_{i=1}^{N(\epsilon)}{\theta^{n}(C_{\rm Smp}(\Gamma_{i}^{\epsilon}))}$ $=$ $\sum_{j=1}^{m}{k_{X,\infty}(C_{j})}\theta^{n}(C_{j}).$
Therefore, from Eq. (1), (2) and
$\theta^{n}(X)=\sum_{i=1}^{N(\epsilon)}{\theta^{n}(\Gamma_{i}^{\epsilon})},$
we obtain that
$\theta^{n}(X)=\sum\limits_{j=1}^{m}k_{X,\infty}(C_{j})\cdot\theta^{n}(C_{j})=\sum\limits_{j=1}^{m}k_{X,\infty}(C_{j})\cdot\frac{\mathcal{H}^{n}(C_{j}\cap B^{{n+k}})}{\mu_{n}}.$
3.2. Some examples
Example 3.2.
Let
$X=\left\{(x,y)\in\mathbb{R}^{2};y-x^{2}=0\right\},$
be an algebraic sets. Then $\theta(X)=1$.
Example 3.3.
Let
$C=\{(x,y,z)\in\mathbb{R}^{3};x^{2}+y^{2}=\cosh^{2}(z)\},$
be the catenoid surface. Then $\theta^{2}(C)=2$.
Figure 1. The catenoid minimal surface (see [6])
Example 3.4.
Let $X=\{(x,y,z)\in\mathbb{R}^{3};z^{2}=\alpha(x^{2}+y^{2})\}$, for $\alpha>0$. Then $\theta(X)=2(1+\alpha)^{-\frac{1}{2}}$ (see [23, Example 2.5]).
Note that if we remove the hypothesis $X$ is a definable set, the density can be infinite or can not even exist.
Example 3.5.
Let
$H=\left\{(x,y,z)\in\mathbb{R}^{3};z=\tan^{-1}\left(\frac{y}{x}\right)\right\}$
be the helicoid surface. Then $\theta^{2}(H)=+\infty$.
Figure 2. The helicoid minimal surface (see [6])
Example 3.6.
Let $I_{1}=(0,a_{1})$ be an open interval where $a_{1}>0$, and we define, by recurrence, the following sets: $A_{1}:=I_{1}\times\{0\}$, $A_{2}:=I_{2}\times\left(\{-\frac{1}{2}\}\cup\{\frac{1}{2}\}\right)$, where $I_{2}=(a_{1},a_{1}+2\mathcal{H}^{1}(I_{1}))$, $A_{3}:=I_{3}\times\{0\}$, where $I_{3}=(a_{2},a_{2}+2\mathcal{H}^{1}(I_{2}))$ and in several cases
$A_{2j+1}=I_{2j}\times\{0\},$
and
$A_{2j}=I_{2j-1}\times\left(\left\{-\frac{1}{2}\right\}\cup\left\{\frac{1}{2}\right\}\right),$
where and $I_{j}=(a_{j-1},a_{j-1}+2\mathcal{H}^{1}(I_{j-1})),$ and $\mathcal{H}^{1}(I_{k})$ is the length of the interval $I_{k}$. Thus we define
$X=\bigcup_{j\in\mathbb{N}}{A_{j}}.$
Then, the following limit does not exist
$\lim_{r\rightarrow+\infty}\frac{\mathcal{H}^{1}(X\cap B_{r}^{2})}{\mu_{1}r^{2}}.$
3.3. First consequences
3.3.1. The degree of an algebraic set is the density at infinity
Corollary 3.7.
Let $X\subset\mathbb{C}^{n}$ be a pure dimensional complex algebraic set with $\dim_{\mathbb{C}}X=d$. Then $\theta^{2d}(X)={\rm deg}(X).$
Proof.
Note that when $X$ is a complex algebraic set, there is a complex analytic set $\sigma$ with $\dim\sigma<\dim X$, such that for each irreducible component $X_{j}$ of tangent cone $C(X,\infty)$, $X_{j}\setminus\sigma$ intersect only one connected component $C_{i}$ of $C_{\rm Smp}(X,\infty)$, then we define $k_{X,\infty}(X_{j}):=k_{X,\infty}(C_{i})$. Thus, from [16, §2.3], we obtain
${\rm deg}(X)=\sum\limits_{j=0}^{r}k_{X,\infty}(X_{j})\cdot{\rm deg}(X_{j}),$
where $X_{1},...,X_{r}$ are all the irreducible components of $C(X,\infty)$. Since each $X_{j}$ is a complex algebraic cone, we have that ${\rm deg}(X_{j})=m(X_{j},0)$, where $m(X_{j},0)$ denotes the multiplicity of $X_{j}$ at 00 (see the definitions of multiplicity and degree in [8]). By [11, Theorem 7.3], $\theta^{2d}(X_{j},0)=m(X_{j},0)$ for each $j\in\{1,...,r\}$. Since each $X_{j}$ is a cone with the vertex at 00, we have that $\theta^{2d}(X_{j},0)=\theta^{2d}(X_{j})$. Thus
${\rm deg}(X)=\sum\limits_{j=0}^{r}k_{X,\infty}(X_{j})\cdot\theta^{2d}(X_{j}).$
By Theorem 3.1, we have that
$\theta^{2d}(X)=\sum\limits_{j=0}^{r}k_{X,\infty}(X_{j})\cdot\theta^{2d}(X_{j}).$
Therefore $\theta^{2d}(X)={\rm deg}(X)$. ∎
3.3.2. The o-minimal Chow’s theorem
Let us remember the Stoll-Bishop’s Theorem:
Theorem 3.8 (Stoll-Bishop’s Theorem (see [45, 46, 5])).
Let $Z\subset\mathbb{C}^{k}$ be a pure dimensional entire complex analytic subset with $\dim_{\mathbb{C}}X=d$. Then $Z$ is a complex algebraic set if and only if there exists a constant $R>0$ such that
$\frac{\mathcal{H}^{2d}(Z\cap\overline{B}_{r}^{2k}(0))}{r^{2d}}\leq R,\quad\mbox{ for all }r>0.$
Thus, we obtain the o-minimal Chow’s Theorem proved in [31, Corollary 4.5]:
Corollary 3.9 (O-minimal Chow’s Theorem).
Let $X\subset\mathbb{C}^{n}$ be a pure dimensional entire complex analytic set with $\dim_{\mathbb{C}}X=d$. Then, $X$ is a complex algebraic set if and only if $X$ is definable in an o-minimal structure on $\mathbb{R}$.
Proof.
By Theorem 3.1, $\theta^{2d}(X)$ is finite. Then there exists a constant $R>0$ such that
$\frac{\mathcal{H}^{2d}(X\cap\overline{B}_{r}^{2k}(0))}{r^{2d}}\leq R,\quad\mbox{ for all }r>0.$
Therefore, by Stoll-Bishop’s Theorem, $X$ is a complex algebraic set. ∎
3.3.3. The Lipschitz Chow’s theorem
Corollary 3.10.
Let $X\subset\mathbb{C}^{n}$ be a pure dimensional entire complex analytic set with $\dim_{\mathbb{C}}X=d$. Then, $X$ is a complex algebraic set if and only if $X$ is lipeomorphic at infinity to a definable set in an o-minimal structure on $\mathbb{R}$.
Proof.
We follow closely the proof of Theorem 3.1 in [38].
We may assume that $0\in X$.
We assume that $X$ is lipeomorphic at infinity to a definable set in an o-minimal structure on $\mathbb{R}$. Let $A\subset\mathbb{C}^{m}$ be a definable set such that there exist compact subsets $K\subset\mathbb{C}^{n}$ and $\tilde{K}\subset\mathbb{C}^{m}$ and a bi-Lipschitz homeomorphism $\varphi\colon A\setminus\tilde{K}\to X\setminus K$. Let $\lambda\geq 1$ such that
$\frac{1}{\lambda}\|x-y\|\leq\|\varphi(x)-\varphi(y)\|\leq\lambda\|x-y\|,\quad\forall x,y\in A\setminus\tilde{K}.$
Fix $x_{0}\in A\setminus\tilde{K}$ and set $y_{0}=\varphi(x_{0})$. Let $\tilde{r}_{0}=\|x_{0}\|$ and $r_{0}=\|y_{0}\|$. Thus, for any $r>0$, we obtain that
$(X\setminus K)\cap B_{r}^{2n}(0)\subset\varphi((A\setminus\tilde{K})\cap B_{\lambda(r+r_{0})+\tilde{r}_{0}}^{2m}(0))$
and
$\mathcal{H}^{2d}(X\cap B_{r}^{2n}(0))$$\leq$ $\lambda^{2d}\mathcal{H}^{2d}(A\cap B_{\lambda(r+r_{0})+\tilde{r}_{0}}^{2m}(0))+\mathcal{H}^{2d}(X\cap K).$
Since $A$ is a definable set, by Theorem 3.1, $\theta^{2d}(A)$ is finite. Then there exists a constant $C>0$ such that
$\frac{\mathcal{H}^{2d}(A\cap\overline{B}_{r}^{2k}(0))}{r^{2d}}\leq C,\quad\mbox{ for all }r>0.$
But $X$ is analytic at 00, there exist $r_{1},R_{1}>0$ such that
$\frac{\mathcal{H}^{2d}(X\cap B_{r}^{2n}(0))}{r^{2d}}\leq R_{1},$
for all $r\leq r_{1}$. Moreover, $\mathcal{H}^{2d}(X\cap K)<+\infty$. Then, there exists $R_{2}>0$ such that
$\frac{\mathcal{H}^{2d}(X\cap B_{r}^{2n}(0))}{r^{2d}}\leq\frac{\lambda^{2d}C(\lambda(r+r_{0})+\tilde{r}_{0})^{2d}+\mathcal{H}^{2d}(X\cap K)}{r^{2d}}\leq R_{2},$
for all $r\geq r_{1}$.
Therefore, by Stoll-Bishop’s Theorem, $X$ is a complex algebraic set. ∎
4. Parametric versions of the Bernstein Theorem
Theorem 4.1.
Let $X\subset\mathbb{R}^{n+k}$ be a closed and connected set and $n=\dim_{H}X$. Assume that $X$ satisfies the monotonicity formula at $p\in X$. Then the following statements are equivalent:
  • (1)
    $X$ is an affine linear subspace;
  • (2)
    $X$ is a definable set, blow-spherical regular at infinity and $C(X,\infty)$ is a linear subspace;
  • (3)
    $X$ is a definable set that is Lipschitz regular at infinity and $C(X,\infty)$ is a linear subspace;
  • (4)
    $X$ is an LNE at infinity definable set and $C(X,\infty)$ is a linear subspace;
  • (5)
    $\theta^{n}(X)=1$.
Proof.
It is obvious that item $(1)$ implies the items $(2)$, $(3)$, $(4)$ and $(5)$.
$(2)\Longrightarrow(5)$. Since $X$ is blow-spherical regular at infinity, by Proposition 2.14 in [40], we have $k_{X,\infty}(C)=1$ for all connected component $C$ of $C_{{\rm Smp}}(X,\infty)$. It follows from the hypothesis that $C(X,\infty)$ is a hyperplane and Theorem 3.1 that $\theta^{d}(X)=1$.
We are going to prove that $(3)\Longrightarrow(4)$.
$(4)\Longrightarrow(5)$. Indeed, since $X$ is a definable set in an o-minimal structure, we have the following:
Claim 4.1.1.
If $X\subset\mathbb{R}^{n}$ is a definable set and LNE at infinity, then $k_{X,\infty}(v)=1$ for all $v\in C_{\rm Smp}(X,{\infty})$.
Proof of Claim 4.1.1.
Indeed, we suppose there exists $x=(x^{\prime},0)\in Smp(\partial X_{\infty}^{\prime})=C(X,\infty)\cap\mathbb{S}^{n-1}\times\{0\}$ such that $k_{X,\infty}(x^{\prime})\geq 2$, we know that $k_{X,\infty}(X_{j})=k_{X,\infty}(x^{\prime})$ is the number of connected components of the germ $(\rho_{\infty}^{-1}(X),x)$, where $X_{j}$ is a connected component of $C_{\rm Smp}(X,{\infty})$ and $x^{\prime}\in X_{j}$. Thus, for a sequence of positive real numbers $\{t_{k}\}_{k\in\mathbb{N}}$ such that $\lim t_{k}=+\infty$, $k_{X,\infty}(x^{\prime})$ can be seen as the number of connected components of the set $(SX\cap\mathbb{S}^{n-1}\times\{t_{k}^{-1}\})\cap B_{\delta}^{n+1}(x),$ for all large $k$ and a small enough $\delta>0$, where
$SX=\{(x,s)\in\mathbb{S}^{n-1}\times(0,+\infty);\,\textstyle\frac{1}{s}\cdot x\in X\mbox{ and }\frac{1}{s}\in S\}$
and $S=\{t_{k}^{-1}\}_{k\in\mathbb{N}}$.
Let $X_{1}$ and $X_{2}$ be two connected components of $X_{\infty}^{\prime}\cap B_{\delta}^{n+1}(x),$ and for $\eta,R>0$, we define the following set $C_{\eta,R}(x^{\prime})=\{v\in\mathbb{R}^{n}\setminus\{0\};\exists t>0;\|v-tx^{\prime}\|\leq\eta t\}\setminus B_{R}^{n}$. For each $j\in\mathbb{N}$, we can take $x_{j}\in\rho_{\infty}(X_{1})$ and $y_{j}\in\rho_{\infty}(X_{2})$ such that $\|x_{j}\|=\|y_{j}\|=t_{j}$ and $\lim_{j\rightarrow+\infty}{\frac{x_{j}}{t_{j}}}=\lim_{j\rightarrow+\infty}{\frac{y_{j}}{t_{j}}}=x^{\prime}$. Now, taking subsequence, if necessary, we assume that $x_{j},y_{j}\in C_{\frac{\eta}{2},R}$. Thus, let $\beta_{j}\colon[0,1]\rightarrow X$ be a curve connecting $x_{j}$ to $y_{j}$, we have that there exists a $t_{0}\in[0,1]$ such that $\beta_{j}(t_{0})\notin C_{\eta,R}$, since $x_{j}$ and $y_{j}$ belong to different connected components of $X\cap C_{\eta,R}$. Therefore, $lenght(\beta_{j})\geq\eta t_{j}$ and then $d_{X}(x_{j},y_{j})\geq\frac{\eta}{2}t_{j}$.
On the other hand, $X$ is a Lipschitz normally embedded at infinity set, then there exists a ball $B_{r}^{n}$ and a constant $C>0$ such that $d_{X\setminus B_{r}^{n}}(v,w)\leq C\|v-w\|$, for all $v,w\in X\setminus B_{r}(0)$. Therefore, $C\|\frac{x_{j}}{t_{j}}-\frac{y_{j}}{t_{j}}\|\geq\eta$, for all $j\in\mathbb{N}$ and this is a contradiction, since $\lim{\frac{x_{j}}{t_{j}}}=\lim{\frac{y_{j}}{t_{j}}}$.
Therefore $k_{X,\infty}(v)=1$ for all $v\in C_{\rm Smp}(X,{\infty})$. ∎
So, since $C(X,\infty)$ is a linear subspace, it follows from Claim 4.1.1, Remark 2.13, and Theorem 3.1 that $\theta^{d}(X)=\mathcal{H}^{d}(C_{1})=1$.
$(5)\Longrightarrow(1)$. Suppose that $\theta^{d}(X)=1$.
Since $X$ satisfies the monotonicity formula at $p$, $\theta^{d}(X,p,\cdot)$ is a non-decreasing function such that $11\leq\theta^{d}(X,p,r)\leq\theta^{d}(X)=$ 1 for all $r>0$. Then $\theta^{d}(X,p,r)=1$ for all $r>0$, and by using again that $X$ satisfies the monotonicity formula at $p$, we have that $X$ is a cone with vertex at $p$ and $X$ is $C^{1}$-smooth at $p$. Therefore $X$ is an affine linear subspace. ∎
4.1. Some examples
Example 4.2.
The catenoid is the surface $C=\{(x,y,z)\in\mathbb{R}^{3};x^{2}+y^{2}=\cosh^{2}(z)\}$. The Catenoid is a complete embedded minimal hypersurface and its geometric tangent cone at infinity is the plane $z=0$. Moreover, it is definable in the o-minimal structure $\mathbb{R}_{exp}$ (see [12]).
Example 4.3.
Let $L\subset\mathbb{S}^{2}$ be a curve lipeomorphic to $\mathbb{S}^{1}$ with $2\pi<\mathcal{H}^{1}(L)<+\infty$. Thus, the surface $X=Cone_{0}(L):=\{tx;x\in L$ and $t\geq 0\}$ is lipeomorphic to $\mathbb{R}^{2}$ and, in particular, it is Lipschitz regular at infinity. Since $X$ is a cone, then $\theta^{2}(X,0,r)=\mathcal{H}^{1}(L)/(2\pi)>1$ for all $r>0$. Therefore $X$ satisfies the monotonicity formula at 00. However, it is not a plane.
Example 4.4.
The surface $P=\{(z,w)\in\mathbb{C}^{2};z=w^{2}\}$ is a smooth complex algebraic curve and, in particular, it is an area-minimizing set. Its geometric tangent cone at infinity is the complex line $w=0$. However, $P$ is not a plane.
Example 4.5.
We cannot drop the hypothesis that $X$ satisfies the monotonicity formula at $p$. Indeed, $X=\{(x,y,z)\in\mathbb{R}^{3};z^{3}=x^{2}+y^{2}+1\}$ is lipeomorphic to $\mathbb{R}^{2}$ and, in particular, it is Lipschitz regular at infinity. In fact, $X$ is the graph of the Lipschitz function $f\colon\mathbb{R}^{2}\to\mathbb{R}$ given by $f(x,y)=(x^{2}+y^{2}+1)^{\frac{1}{3}}$. Moreover, $X$ is a smooth surface, $C(X,\infty)$ is the plane $z=0$ and $\theta^{2}(X)=1$. However, $X$ is not a plane.
4.2. Some direct consequences
It is a direct consequence of Theorem 3.1 and Claim 4.1.1 the following result:
Corollary 4.6.
Let $X\subset\mathbb{R}^{n+d}$ be an $n$-dimensional definable set. If $X$ is LNE at infinity, then $\theta^{n}(X)=\theta^{n}(C(X,\infty))$.
In particular, since for a pure dimensional complex algebraic set, $X\subset\mathbb{C}^{n+d}$ with $\dim_{\mathbb{C}}X=d$, we have $\theta^{2d}(X)={\rm deg}(X)$ (see Corollary 3.7), we obtain the main result of [10]:
Corollary 4.7.
Let $X\subset\mathbb{C}^{n+d}$ be pure dimensional algebraic set. If $X$ is LNE at infinity, then $\deg(X)=\deg(C(X,\infty))$.
Corollary 4.8.
Let $X\subset\mathbb{R}^{n+k}$ be a pure dimensional algebraic set. Suppose that $X$ is LNE at infinity and $C(X,\infty)$ is a linear subspace. If $X$ is a minimal submanifold or an area-minimizing set, then $X$ is an affine linear subspace.
Proof.
By Corollary 4.6, $\theta^{d}(X)=\theta^{d}(C(X,\infty))$. Since $C(X,\infty)$ is a hyperplane, we have $\theta^{d}(X)=1$. By Theorem 4.1, $X$ is an affine linear subspace. ∎
We obtain the main results of [14] (see also [38]):
Corollary 4.9.
Let $X\subset\mathbb{C}^{n+k}$ be a pure dimensional complex analytic set with $\dim_{\mathbb{C}}X=d$. Suppose that $X$ is LNE at infinity and $C(X,\infty)$ is a dimensional complex linear subspace with $\dim_{\mathbb{C}}C(X,\infty)=d$. Then $X$ is an affine linear subspace.
Proof.
Since $C(X,\infty)$ is a dimensional complex algebraic set with $\dim_{\mathbb{C}}C(X,\infty)=d$, we have that $X$ is a complex algebraic set as well (see, e.g., [40, Theorem 3.1]). In particular, $X$ is a definable set and satisfies the monotonicity formula at any $p\in X$. By Theorem 3.1, $X$ is an affine linear subspace. ∎
Corollary 4.10.
Let $X\subset\mathbb{C}^{n+d}$ be a pure dimensional complex algebraic set with $\dim_{\mathbb{C}}X=d$. Suppose that $X$ is Lipschitz regular at infinity. Then $X$ is an affine linear subspace.
Proof.
$X$ is Lipschitz regular at infinity, then $X$ is LNE at infinity, and by Theorem 2.12, $C(X,\infty)$ is lipeomorphic to $\mathbb{R}^{2d}$. In particular, $C(X,\infty)$ is topological manifold, and thus by Prill’s theorem ([33, Theorem]), $C(X,\infty)$ is a dimensional complex linear subspace with $\dim_{\mathbb{C}}C(X,\infty)=d$. By Theorem 3.1, $X$ is an affine linear subspace. ∎
We obtain also the following version of Moser’s result [27, Corollary, p. 591]:
Corollary 4.11.
Let $X\subset\mathbb{R}^{n+1}$ be a complete minimal hypersurface. Suppose that there are compact sets $K\subset\mathbb{R}^{n}$ and $\tilde{K}\subset\mathbb{R}^{n+1}$ such that $X\setminus\tilde{K}$ is the graph of a definable Lipschitz function $u\colon\mathbb{R}^{n}\setminus K\to\mathbb{R}$. Then $u$ is the restriction of an affine function and, in particular, $X$ is a hyperplane.
Proof.
We need a preliminary result. For a set $A\subset\mathbb{R}^{n}$, denote by $\mathcal{N}(A,\infty)$ the union of all hyperplanes $T$ such that there is a sequence $\{x_{j}\}_{j}\subset A\setminus{\rm Sing}_{1}(A)$ such that $\lim\|x_{j}\|=+\infty$ and $T_{x_{j}}A$ converges to $T$.
Lemma 4.12.
Let $A\subset\mathbb{R}^{n}$ be an unbounded definable set. Then $C(A,\infty)\subset\mathcal{N}(A,\infty)$.
Proof.
Firstly, note that $C(A\setminus{\rm Sing}_{1}(A),\infty)=C(A,\infty)$ and $\mathcal{N}(A\setminus{\rm Sing}_{1}(x),\infty)=\mathcal{N}(A,\infty)$. Thus, we may assume that $A$ is $C^{1}$ smooth.
Let $v\in C(A,\infty)$. We are going to show that $v\in\mathcal{N}(A,\infty)$ and we may assume that $\|v\|=1$. In order to do that, it is enough to find, for some $R>0$, a definable $C^{1}$ smooth arc $\alpha\colon(R,+\infty)\to A$ such that $\alpha(t)=tv+o_{\infty}(t)$. Indeed, by taking a subsequence, if necessary, we may assume that $T_{\alpha(j)}A$ converges to $T\subset\mathcal{N}(A,\infty)$. Since $\alpha^{\prime}(j)\in T_{\alpha(j)}A$ for all $j$, then $\lim\alpha^{\prime}(j)=v\in T$.
Now, let us prove that there is such an arc $\alpha$. Let $Y=\rho_{\infty}^{-1}(A)$. We have that $Y$ is a definable set. Note that the closure of $Y$ in $\mathbb{S}^{n-1}\times[0,\infty)$ satisfies the following:
$\overline{Y}\cap(\mathbb{S}^{n-1}\times\{0\})=(C(A,\infty)\cap\mathbb{S}^{n-1})\times\{0\}.$
Thus, $(v,0)\in\overline{Y}$. By the Curve Selection Lemma, there is a definable continuous arc $\beta\colon[0,\epsilon)\to\overline{Y}$ such that $\beta((0,\epsilon))\subset Y$ and $\beta(0)=(v,0)$. Let us write $\beta(s)=(x(s),h(s))$. Since $h$ is not a constant function around 00, by Monotonicity Lemma, one can suppose that $\beta$ is $C^{1}$ smooth and that $h$ is $C^{1}$ smooth, strictly increasing in the domain $(0,\delta)$ and $h^{\prime}(s)\not=0$ for all $s\in(0,\delta)$. Hence, $h\colon(0,\delta/2)\to(0,\delta^{\prime})$ is a $C^{1}$ diffeomorphism, where $\delta^{\prime}=h(\frac{\delta}{2})$. Let $\alpha\colon(\frac{1}{\delta^{\prime}},+\infty)\to A$ be the arc $\alpha(t)=\rho_{\infty}\circ\beta\circ h^{-1}(\frac{1}{t})$.
Clearly, $\alpha$ is $C^{1}$ smooth, and since $\lim\limits_{s\to 0^{+}}x(s)=x(0)=v$, we obtain
$\alpha(t)=tx(h^{-1}(\frac{1}{t}))=tv+o_{\infty}(t).$
From the above lemma, we obtain that $C(X,\infty)\subset\mathcal{N}(X,\infty)$.
By Lemma 1.18 in [43], we have that the following limit exists:
$\lim\limits_{x\to\infty}(Du(x),-1)=w.$
Therefore, $\mathcal{N}(X,\infty)=w^{\perp}$. In particular, $C(X,\infty)$ is a subset of an $n$-dimensional hyperplane, and thus $\mathcal{H}^{n}(C(X,\infty)\cap B_{1}^{n+1})\leq\mu_{n}$. Since there is a blow-spherical homeomorphism at infinity between $X$ and $\mathbb{R}^{n}$, then there is only one relative multiplicity which is $1$. Then, by Theorem 3.1, $\theta^{n}(X)\leq 1$. However, since $X$ is a minimal submanifold, $\theta^{n}(X)\geq 1.$ Therefore $\theta^{n}(X)=1$ and by Theorem 4.1, $X$ is a hyperplane. It is easy to verify that $u$ is the restriction of an affine function. ∎
In the next section, we generalize Corollary 4.11 (see Corollary 5.6).
5. Generalization of the Moser’s Bernstein Theorem
In this section, we generalize the Moser’s Bernstein Theorem.
Theorem 5.1.
Let $X\subset\mathbb{R}^{n+k}$ be a connected closed set with Hausdorff dimension $n$. Then, $X$ is a $n$-dimensional affine linear subspace if and only if we have the following:
  1. (1)
    $X$ satisfies the monotonicity formula at $p\in X$;
  2. (2)
    there are compact sets $K\subset\mathbb{R}^{n}$ and $\tilde{K}\subset\mathbb{R}^{n+k}$ such that $X\setminus\tilde{K}$ is the graph of a $C^{1}$ smooth function $u\colon\mathbb{R}^{n}\setminus K\to\mathbb{R}^{k}$ with bounded derivative;
  3. (3)
    $\mathcal{N}(X,\infty)$ is a linear subspace.
Proof.
It is obvious that if $X$ is a $n$-dimensional affine linear subspace, then $X$ satisfies (1), (2) and (3).
Reciprocally, assume that $X$ satisfies (1), (2) and (3).
As before, for a set $A\subset\mathbb{R}^{n}$, denote by $\mathcal{N}(A,\infty)$ the union of all hyperplanes $T$ such that there is a sequence $\{x_{j}\}_{j}\subset A\setminus{\rm Sing}_{1}(A)$ such that $\lim\|x_{j}\|=+\infty$ and $T_{x_{j}}A$ converges to $T$.
Let $P$ be the hyperplane $\mathcal{N}(X,\infty)$. We choose linear coordinates $(y_{1},...,y_{n+k})$ of $\mathbb{R}^{n+k}$ such that $P=\{(y_{1},...,y_{n+k})\in\mathbb{R}^{n+k};y_{n+1}=...=y_{n+k}=0\}$. For a larger enough $R>0$, $X\setminus(\overline{B_{R}^{n}}\times\mathbb{R}^{k})$ is the graph of a function $v=(v_{1},...,v_{k})\colon P\setminus\overline{B_{R}^{n}}\to\mathbb{R}^{k}$. By the Implicit Mapping Theorem, we have that $v$ is a Lipschitz function and that the following limit exists:
$\lim\limits_{y\to\infty}Dv_{i}(y)=\tilde{w}_{i},\quad i\in\{1,...,k\}.$
Therefore $\tilde{w}=(\tilde{w}_{1},...,\tilde{w}_{k})=0$. Then, for each $\varepsilon>0$, there is $r>0$ such that $\|Dv(y)\|\leq\varepsilon$ for all $y\in\Omega:=P\setminus\overline{B_{r}^{n}}$.
Claim 5.1.1.
$v|_{\Omega}$ is a Lipschitz mapping with constant $\pi\varepsilon$.
Proof.
Firstly, we define the set of all piecewise $C^{1}$-smooth curves connecting $x$ to $y$ in $\Omega$. We denote such a set by:
$\Omega(x,y):=\left\{\gamma\colon[0,1]\rightarrow\Omega;\gamma(0)=x,\gamma(1)=y\mbox{ and }\gamma\mbox{ is piecewise $C^{1}$ smooth}\right\}.$
Let $x,y\in\Omega$ be any two points. Then
[…]
$=$ $\varepsilon d_{\Omega}(x,y).$
However, $\Omega$ is an LNE set such that $d_{\Omega}(x,y)\leq\pi\|x-y\|$ for all $x,y\in\Omega$. Therefore,
$\|v(x)-v(y)\|\leq\pi\varepsilon\|x-y\|.$
Claim 5.1.2.
$1\leq\theta^{n}(X)\leq(1+\pi\varepsilon)^{n}.$
Proof.
Let $\varphi\colon\Omega\rightarrow A=X\setminus(\overline{B_{r}^{n}}\times\mathbb{R}^{k})$ be the mapping given by $\varphi(x)=(x,v(x))$. We have that $\varphi$ is a bi-Lipschitz mapping such that
$\|x-y\|\leq\|\varphi(x)-\varphi(y)\|\leq(1+\pi\varepsilon)\|x-y\|.$
Therefore,
$\mathcal{H}^{n}(B^{n}_{t}\setminus B^{n}_{r})\leq\mathcal{H}^{n}(A\cap B_{t}^{n+k})\leq(1+\pi\varepsilon)^{n}\mathcal{H}^{n}(B^{n}_{t+\pi\varepsilon}),$
for all $t>r$. Thus, we obtain the following $\theta^{n}(\mathbb{R}^{n}\setminus B^{n}_{r})\leq\theta^{n}(X)\leq(1+\pi\varepsilon)^{n}\theta^{n}(\mathbb{R}^{n}).$ Since $\theta^{n}(\mathbb{R}^{n}\setminus B^{n}_{r})=\theta^{n}(\mathbb{R}^{n})=1$ and $\theta^{n}(A)=\theta^{n}(X)$, we obtain that $1\leq\theta^{n}(X)\leq(1+\pi\varepsilon)^{n}$, which finishes the proof of the claim. ∎
Setting $\varepsilon\to 0^{+}$, we obtain that $\theta^{n}(X)=1$. By Theorem 4.1, $X$ is an affine linear subspace.
5.1. Examples
In this subsection, by presenting several examples, we show that Theorem 5.1 is sharp in the sense that no hypothesis of it can be dropped.
We cannot remove the hypothesis that $\mathcal{N}(X,\infty)$ is a linear subspace.
Example 5.2 (Theorem 7.1 in [24]).
There is a connected closed semialgebraic set $X\subset\mathbb{R}^{n+k}$ which is an area-minimizing set and is the graph of a semialgebraic Lipschitz function $u\colon\mathbb{R}^{n}\to\mathbb{R}^{k}$. Moreover, $X$ is not an affine linear subspace.
We cannot remove the condition that $X$ is the graph of a function outside of a compact.
Example 5.3.
The catenoid $C=\{(x,y,z)\in\mathbb{R}^{3};x^{2}+y^{2}=\cosh^{2}(z)\}$ (see Example 4.2) satisfies the monotonicity formula at any point $p\in C$ and $N(C,\infty)$ is the plane $z=0$, but $C$ is not a plane.
We cannot drop the hypothesis that the function $u$ has bounded derivative.
Example 5.4.
The surface $P=\{(z,w)\in\mathbb{C}^{2};z=w^{2}\}$ is a smooth complex algebraic curve and, in particular, it is an area-minimizing set. $N(P,\infty)$ is the complex line $w=0$. However, $P$ is not a plane.
We cannot drop the hypothesis that $X$ satisfies the monotonicity formula at $p$.
Example 5.5.
$X=\{(x,y,z)\in\mathbb{R}^{3};z^{3}=x^{2}+y^{2}+1\}$ is the graph of the smooth Lipschitz function $f\colon\mathbb{R}^{2}\to\mathbb{R}$ given by $f(x,y)=(x^{2}+y^{2}+1)^{\frac{1}{3}}$. Moreover, $X$ is a smooth surface and $\mathcal{N}(X,\infty)$ is the plane $z=0$. However, $X$ is not a plane.
5.2. Consequences
Corollary 5.6.
Let $X\subset\mathbb{R}^{n+1}$ be a connected closed set with Hausdorff dimension $n$ and that satisfies the monotonicity formula at $p\in X$. Suppose that there are compact sets $K\subset\mathbb{R}^{n}$ and $\tilde{K}\subset\mathbb{R}^{n+1}$ such that $X\setminus\tilde{K}$ is a minimal hypersurface that is the graph of a $C^{2}$ smooth function $u\colon\mathbb{R}^{n}\setminus K\to\mathbb{R}$ with bounded derivative. Then $u$ is the restriction of an affine function and, in particular, $X$ is a hyperplane.
Proof.
By Lemma 1.18 in [43] and Theorem X in [2], we have that the following limit exists:
$\lim\limits_{x\to\infty}(Du(x),-1)=w.$
Therefore, $\mathcal{N}(X,\infty)$ is the hyperplane $w^{\perp}$.
By Theorem 5.1, $X$ is a hyperplane. It is obvious that $u$ is the restriction of an affine function. ∎
Corollary 5.7.
Let $X\subset\mathbb{R}^{n+1}$ be a connected closed set with Hausdorff dimension $n$ and that satisfies the monotonicity formula at $p\in X$. Suppose that $2\leq n\leq 7$ and there are compact sets $K\subset\mathbb{R}^{n}$ and $\tilde{K}\subset\mathbb{R}^{n+1}$ such that $X\setminus\tilde{K}$ is a minimal hypersurface that is the graph of a $C^{2}$ smooth function $u\colon\mathbb{R}^{n}\setminus K\to\mathbb{R}$. Then $u$ is the restriction of an affine function and, in particular, $X$ is a hyperplane.
Proof.
By Theorem 1 in [43] and Theorem X in [2], $Du$ is bounded. By Corollary 5.6, $X$ is a hyperplane and $u$ is the restriction of an affine function. ∎
Since a minimal hypersurface satisfies the monotonicity formula at any of its points, we obtain, as direct consequences of Corollaries 5.7 and 5.6, the following results:
Corollary 5.8.
Let $X\subset\mathbb{R}^{n+1}$ be a connected complete minimal hypersurface. Suppose that there are compact sets $K\subset\mathbb{R}^{n}$ and $\tilde{K}\subset\mathbb{R}^{n+1}$ such that $X\setminus\tilde{K}$ is the graph of a $C^{2}$ smooth function $u\colon\mathbb{R}^{n}\setminus K\to\mathbb{R}$ with bounded derivative. Then $X$ is a hyperplane and $u$ is the restriction of an affine function.
Corollary 5.9.
Let $X\subset\mathbb{R}^{n+1}$ be a connected complete minimal hypersurface with $2\leq n\leq 7$. Suppose that there are compact sets $K\subset\mathbb{R}^{n}$ and $\tilde{K}\subset\mathbb{R}^{n+1}$ such that $X\setminus\tilde{K}$ is the graph of a $C^{2}$ smooth function $u\colon\mathbb{R}^{n}\setminus K\to\mathbb{R}$. Then $X$ is a hyperplane and $u$ is the restriction of an affine function.
We finish this section by presenting the following example, which shows that the hypothesis that $X$ is a complete hypersurface in the above corollaries cannot be dropped:
Example 5.10.
[…]
$\left|\frac{\partial f}{\partial x}(x,y)\right|$ $=$ $\frac{x}{\sqrt{x^{2}+y^{2}}}\frac{1}{\sqrt{x^{2}+y^{2}-1}}\leq 1,$
and
$\left|\frac{\partial f}{\partial y}(x,y)\right|$ $=$ $\frac{y}{\sqrt{x^{2}+y^{2}}}\frac{1}{\sqrt{x^{2}+y^{2}-1}}\leq 1.$
Since $\mathbb{R}^{2}\setminus\overline{B_{2}^{2}}$ is Lipschitz normally embedded set, we have that $f$ is Lipschitz (see Claim 5.1.1).
Acknowledgements. The authors would like to thank Alexandre Fernandes and Pacelli Bessa for their interest, incentive and comments on this research.